OpenStudy (anonymous):

Five charges are placed in a closed box. Each charge (except the first) has a magnitude which is twice that of the previous one placed in the box. All charges have the same sign and (after all the charges have been placed in the box) the net electric flux through the box is 9.5 × 10 7 N · m2 /C. What is the magnitude of the smallest charge in the box? The permittivity of a vacuum is 8.85419 × 10 −12 C 2 /N · m2 . Answer in units of µC

OpenStudy (jamesj):

Remember the Gaussian result that the integral of the flux must be equal to the sum of the charges \[ \int E \cdot dA = \frac{\sum Q}{\epsilon_0} \] Now you know the left hand side. Use that to solve for the RHS.

OpenStudy (anonymous):

^^ Do what JamesJ said

OpenStudy (turingtest):

yep!

OpenStudy (jamesj):

Notice that the sum of the charges is \[ q + 2q + 4q + 8q + 16q = 31q \]

OpenStudy (anonymous):

so (9.5 * 10^7)(8.85 *10^-12)/31

OpenStudy (jamesj):

Looks about right.

OpenStudy (anonymous):

my answer is 2,713 * 10^-5. imentering this in the system but its coming wrong any ideas why?

OpenStudy (jamesj):

Given the data you provided, that is the answer.

OpenStudy (turingtest):

^agreed